OpenStudy (anonymous):

From a balloon 1191 feet high, the angle of depression to the ranger headquarters is 72°54'. How far is the headquarters from a point on the ground directly below the balloon (to the nearest foot)?

OpenStudy (anonymous):

ok i did 1191 tan(72.9) in degrees and got 3871 but it doesn't match any one of these answers. lost what did i do wrong? 356 ft 361 ft 366 ft 371 ft

OpenStudy (anonymous):

let the required distance be x ft now from the given data x/1191 =cot(72deg54min) hence x=1191*cot(75deg54min) =366 ft(appx) (use calc or log to evaluate )

OpenStudy (anonymous):

lol just found it thx for the help tho